(8%) Problem 5: Suppose a 65 kg gymnast is climbing a rope. atano acctan0 io
Answer #1

Solution:

A) T = mg

=> T = 65 * 9.8

=> T = 637 N

B) T = m(g+a)

=> T = 65(9.8 + 1.4)

=> T = 728 N

thanks